### v2

# A benchmark for computational analysis of animal behavior, using animal-borne tags

Abstract
Animal-borne sensors (‘bio-loggers’) can record a suite of kinematic and environmental data, which can elucidate animal ecophysiology and improve conservation efforts. Machine learning techniques are used for interpreting the large amounts of data recorded by bio-loggers, but there exists no common framework for comparing the different machine learning techniques in this domain. To address this, we present the Bio-logger Ethogram Benchmark (BEBE), a collection of datasets with behavioral annotations, as well as a modeling task and evaluation metrics. BEBE is to date the largest, most taxonomically diverse, publicly available benchmark of this type, and includes 1654 hours of data collected from 149 individuals across nine taxa. In addition, using BEBE, we test a novel self-supervised learning approach to identifying animal behaviors based on bio-logger data, using a deep neural network pre-trained with self-supervision on data collected from human wrist-worn accelerometers. We show that this approach out-performs common alternatives, especially in a setting with a low amount of training data. Datasets, models, and evaluation code are made publicly available at https://github.com/earthspecies/BEBE, to enable community use of BEBE as a point of comparison in methods development.
[…]
Keywords Machine Learning  $\cdot$ Bio-loggers  $\cdot$ Animal Behavior  $\cdot$ Accelerometers  $\cdot$ Time series  $\cdot$ Self-Supervised Learning
Animal behavior is of central interest in ecology and evolution, because an individual’s behavior affects its reproductive opportunities and probability of survival [17]. Additionally, understanding animal behavior can be key to identifying conservation problems and planning successful management interventions [4], for example in rearing captive animals prior to reintroduction [83], designing protected areas [76], and reducing dispersal of introduced species [78].
One increasingly utilized approach for monitoring animal behavior is remote recording by animal-borne tags, or bio-loggers [67, 86, 87]. These tags can be composed of multiple sensors such as accelerometer, gyroscope, altimeter, pressure, GPS receiver, microphone, and camera, which record time-series data on an individual’s behavior and their in situ environment. Additionally, bio-logger datasets can include data from many-hour tag deployments on multiple individuals.
To give a behavioral interpretation to recorded bio-logger data, it is useful to construct an inventory of what types of actions an individual may perform [3]. This inventory, or ethogram, is then used to classify observed actions (Figure 1A). Using an ethogram, one can quantify, for example, the proportion of time an animal spends in different behavioral states, and how these differ between groups (e.g., sex, age, populations), or change over time (e.g., seasonally), with physiological condition (e.g., healthy vs. sick) or across different environmental contexts (e.g. [41]).
For classifying the behaviors underlying bio-logger data, researchers are increasingly using supervised machine learning (ML) techniques [81]. In a typical workflow, a human annotates some of the recorded bio-logger data with the tagged individuals’ behavioral states using a pre-determined ethogram, based on observations made simultaneously with data recording. These annotated data are used to train a ML model, which is then used to predict behavioral labels for the remaining un-annotated portion of the dataset. A test dataset, which is held out from the training stage, can be used to evaluate how well the trained model is able to perform this behavior classification task. In this manner, ML can help scientists to minimize manual effort required to ascribe behavioral labels to bio-logger data, or extend behavioral labels to data where manual ground-truthing is not possible.
Much research has applied ML to bio-logger data to establish its use with particular species, as well as to investigate the impact of different decisions made when using ML models for this purpose. Research questions include characterizing which methods are the most accurate, precise, sensitive, interpretable, or rapid (e.g., [52, 73, 9, 20, 34, 60, 85, 42]), how to reduce the extent of ground-truthing necessary in species that are difficult to observe (e.g., [57, 11, 74, 69, 45, 29, 26]), and what kinds of behaviors are detectable with particular sensors and models (e.g., [40, 73, 75, 25, 70, 16, 51]). However, the majority of studies focus on data from a single or a few closely-related species, making it difficult to identify patterns in how behavior classification methods are applied across multiple datasets.
Figure 1: A) Examples of ethograms in BEBE. Left: gull ethogram with three behaviors. Right: a subset of the dog ethogram, with four behaviors. B) BEBE consists of a supervised behavior classification task on nine annotated datasets, along with a set of metrics that compare model predictions with the annotations. Datasets and code are publicly available at https://github.com/earthspecies/BEBE. C) Datasets in BEBE, with a photo of a representative individual and a 5-minute clip of annotated tri-axial accelerometer (TIA) data for each. Each accelerometer channel is min-max scaled for visualization. Top row: black-tailed gull (Larus crassirostris[39], domestic dog (Canis familiaris[40, 82], carrion crow (Corvus corone[72] (see Methods). Middle row: western diamondback rattlesnake (Crotalus atrox[20], humpback whale (Megaptera novaeangliae[27], New Zealand fur seal (Arctocephalus forsteri[43]. Bottom row: polar bear (Ursus maritimus[56, 57], sea turtle (Chelonia mydas[34], human (Homo sapiens[2]. Gaps indicate that the behavior annotation is Unknown. For image attributions, see acknowledgments.
A commonly used tool in ML for improving our understanding of analysis techniques is the benchmark (e.g. [66]). A benchmark consists of a publicly available dataset, a problem statement specifying a model’s inputs and the desired outputs (a task), and a procedure for quantitatively evaluating a model’s success on the task (using one or several evaluation metrics). In a common use-case for a benchmark, researchers report the performance of a proposed technique on the benchmark, helping the field to draw comparisons between different techniques and consolidate knowledge about promising directions. Developing benchmarks has been identified as an area of focus for ML applications in wildlife conservation [80] and animal behavior [53, 13].
For behavior classification from bio-loggers, a benchmark could assess model performance across a breadth of study systems in order to identify relevant patterns, such as how modeling decisions can influence classification performance. Indeed, researchers have analyzed multi-species datasets to identify best practices for other key challenges in bio-logging, such as sensor calibration [28] and signal processing [46, 62]. To our knowledge, five previous studies have applied one or more behavior classification techniques on multiple bio-logger datasets, and the datasets used in these studies could hypothetically be used as a benchmark [11, 6, 88, 71, 85]. However, four of these do not make all of their bio-logger data publically available[11, 6, 88, 71] and therefore are not available for community use as a benchmark. The fifth does not reflect deployment variability in bio-logger data due to the inclusion of a single individual for each of four species [85].
In order to fill this gap, we present the Bio-logger Ethogram Benchmark (BEBE), designed to capture challenges in behavior classification from diverse bio-logger datasets. BEBE combines nine datasets collected by various research groups, each with behavioral annotations, as well as a supervised behavior classification task with corresponding evaluation metrics (Figure 1B). These datasets are diverse, spanning multiple species, individuals, behavioral states, sampling rates, and sensor types (Figure 1C), as well as large in size, ranging from six to over a thousand hours in duration. We focus on data collected from tri-axial accelerometers (TIA), in addition to gyroscopes, and environmental sensors. TIA are widely incorporated into bio-loggers because they are inexpensive and lightweight [67]. Additionally, the data they collect has been used to infer behavioral states on the order of seconds, in a wide variety of species [86].
As a first application of BEBE, we make and test several hypotheses about ML usage in bio-logger data (Table 1). We base these hypotheses on recent trends in ML, as well as based on their potential to influence the workflow of researchers using ML with bio-logger data. First, in many applications of ML, deep neural networks that make predictions based on raw data, out-perform classical ML methods, which make predictions based on hand-crafted summary statistics or features [44]. In practice, the most commonly used ML methods for bio-logger data have been classical techniques such as random forests [77, Table 2], for which feature engineering is a key challenging step. Contrary to this practice, we predict the dominance of deep neural networks will carry over to bio-logger data (H1).
Second, we examine how data recorded for one species can be used to inform behavior predictions for a different species. Some previous works (e.g. [55, 57, 11]) have adopted a cross-species transfer learning strategy, by training a supervised model on one species and then applying this trained model to another related species. Our approach is different in that we use self-supervised learning. In self-supervised learning, a ML model (typically a deep neural network) is pre-trained to perform an auxiliary task on an unlabeled dataset. Importantly, training the model to perform this auxiliary task does not require any human-generated annotations of the data. In this way, self-supervised pre-training can make use of a large amount of un-annotated data that is easy to obtain (e.g., a species where a large dataset exists). Later, the pre-trained model can be trained (or fine-tuned) to perform the task of interest (such as behavior classification in a different species), using a small amount of annotated data. By learning to perform the auxiliary task the model learns a set of features, which often provide a good set of initial model parameters for performing the task introduced in the fine-tuning step. Inspired by the recent success of self-supervised learning in other domains of ML (e.g. language [10] and computer vision [12]), we predict that a deep neural network pre-trained on a large amount of human accelerometer data will outperform alternative methods, after fine-tuning (H2). Moreover, the self-supervised pre-training step can reduce the amount of annotated data required to meet a given level of performance [14]. Therefore, we predict that this trend will hold when we reduce the amount of training data by a factor of four (reduced data setting; H3).
Finally, we investigate how the performance of our best models varies by behavior class and how this per-behavior performance scales with the amount of training data. In recent years, performance on some human behavior classification benchmarks has shown little improvement in spite of methodological advancements [79]. This suggests that the sensor data may not contain sufficient information to discriminate activities of interest. As found for human activity recognition and behavior classification in other animals [16, 25, 40, 51, 70, 73], we expect that there will be a large degree of variation in per-behavior performance. Here, we test one possibility for improving classification performance: increasing the amount of training data. If some behaviours are not well discriminated by sensor data, we predict that increasing the amount of training data will show only minimal improvement (H4).
While we focus on a specific classification task in this study, all datasets, models, and evaluation code presented in BEBE are available at https://github.com/earthspecies/BEBE for general community use. Researchers may use the standardized task to test classification methods, or adapt BEBE datasets for their own research questions. Given that one aim of BEBE is to improve our understanding of classification methods in bio-logger data, we are also seeking contributions to create an expanded benchmark with improved taxonomic coverage, a broader range of sensor types, additional standardization, and a wider variety of modeling tasks. Details about how to contribute in this way can also be found at our GitHub.
Hypothesis Hypothesis confirmed?
(H1) Deep neural network-based approaches will outperform classical approaches based on hand-chosen summary statistics.
Yes
(H2) Self-supervised pre-training using human accelerometer data will improve classification performance.
Partly
(H3) Self-supervised pre-training using human accelerometer data will improve classification performance when the amount of training data is reduced by a factor of four by removing individuals.
Yes
(H4) In terms of a single model’s predictive performance, there is minimal improvement in some behavior classes when increasing the amount of training data by four times by adding individuals.
Yes
Table 1: Hypotheses tested in this work. BEBE provides a means to identify patterns in behavior classification methods applied across multiple species and sensor types.
1 Results
1.1 Benchmark Datasets
We brought together nine animal motion datasets into a benchmark collection called the Bio-logger Ethogram Benchmark (BEBE) (Table 1.1). BEBE introduces a novel, previously unpublished dataset (Crow); otherwise, these data were all collected in previous studies. Of the datasets included in BEBE, four are publicly available for the first time (Whale, Crow, Rattlesnake, Gull) and five were already publicly available (HAR, Polar bear, Sea turtle, Seals, Dog).
In each dataset, data were recorded by bio-loggers attached to several different individuals of the given species. Each dataset contains one species, except for the Seals dataset which contains four Otariid species. These bio-loggers collected kinematic and environmental time series data, such as acceleration, angular velocity, pressure, and conductivity (Figure 2). While each dataset in BEBE includes acceleration data, different hardware configurations were used across studies. As a result, each dataset comes with its own particular set of data channels, and with its own sampling rate. We used calibrated data as provided by the original dataset authors.
Table 2: Summary of datasets in BEBE. Out of nine datasets, one comes from humans, three come from other terrestrial species, three come from aquatic species, and two come from flying species. Datasets marked with an asterisk are publicly available for the first time in BEBE. Duration: the total duration of the dataset. Duration annotated: the duration of annotated data. Mean annot. dur.: the average duration an individual spends in a behavioral state (described in Methods). Annot. method: the type of data that was used to make the behavioral annotations. $C$: Creative Commons.
HAR
[2, 65]
[…]
[20]
Polar bear
[56, 57]
Dog
[40, 82]
[…]
[27]
Sea Turtle
[34]
Seals
[43, 42]
[…]
[39]
[…]
(see Methods)
Species
Human Western diamondback rattlesnake
Polar bear
Domestic dog Humpback whale
Green turtle
Otariid spp. Black-tailed gull Carrion crow
Typical
Body
mass
60-70 1-3
150-300 (F),
300-800 (M)
5-70 40000
68-190
Fur seals:
40 (F), 140 (M)
Sea lion:
100-300
.4-.6 .4-.6
Location
Genoa,
Italy
Texas, USA Beaufort Sea,
Arctic Ocean
Helsinki,
Finland
Wilhelmina Bay,
Western
Antarctica
Peninsula
Grande Anse
d’Arlet,
Martinique,
France
West coast of
Australia
Kabushima Island,
Japan
León,
Spain
Tag
position
Waist Body
Neck
Back and neck Dorsal surface or flank
Carapace
Back Back or abdomen Tail
# indiv.
30 13
5
45 8
14
12 11 11
Behavior
classes
Lie down
Stand
Sit
Walk downstairs
Walk upstairs
Walk
Moving
Not moving
Walk
Swim
Run
Roll
Rest
Pounce
Head shake
Groom
Eat
Dig
Walk
Trot
Stand
Sniff
Sit
Shake
Pant+stand
Pant+sit
Pant+lie on chest
Lie on
chest
Gallop
Feed dive
Exploratory dive
Travel dive
Rest dive
Swim
Stay at
surface
Scratch
Rest
Glide
Feed
Breathe
Feed
Groom
Rest
Travel
Fly
Forage
Stationary
In nest
Fly
Sample
rate (Hz)
50 1
16
100 5
20
25 25 50
Data
channels
TIA
Gyroscope
TIA TIA
Conductivity
2x TIA
2x gyroscope
TIA
Depth
Speed
TIA
Gyroscope
Depth
TIA
Depth
TIA TIA
Duration (hrs)
6.2 30.9
1108.4
29.5 184.6
77.1
14.0 85.7 114.6
Duration annotated (hrs)
4.2 30.9
196.1
16.9 114.1
67.8
11.6 85.0 3.4
Mean
annot.
dur. (sec)
17.5 721.8
127.2
15.5 119.8
47.2
24.8 2823.7 14.1
Annot.
method
Direct obs. Direct obs.
On tag
video
Off tag
video
Motion On tag
video
Off tag
video
On tag
video
On tag
audio
License
Custom $C$
Public
Domain
$C$ $C$
Public
Domain
$C$ $C$
$C$
In addition to the time series bio-logger data, each dataset in BEBE comes with human-generated behavioral annotations (Figure 2, colored bars). In each dataset, each sampled time step is annotated with the current behavioral state of the tagged individual, which can be one of several discrete behavioral classes. At some time steps, it was not possible to observe the individual, or it was not possible to classify the individual’s behavior using the predefined behavioral classes. In these cases, this time step is annotated as Unknown. These Unknown behavioral annotations are disregarded during model training and evaluation (also see Methods).
There are multiple time scales of behavior represented across the nine ethograms in BEBE, with some datasets including brief activities (e.g. shaking), and some including longer duration activities (e.g. foraging). In Table 1.1 we report the mean duration (in seconds) of an annotation in each dataset, as a rough estimate of the mean duration an individual spends in a given behavioral state.
In addition to Figure 2, we provide additional data visualizations in the Supplementary Information. Examples of each behavior class with the full set of channels reveal varying degrees of stereotypy in the behavioral classes (Supplementary Figures S1-S10). For example, in the Sea Turtle dataset, Stay at surface appears more stereotyped than Feed. Summary statistics across different individuals for each behavior class suggest the presence of discriminative features for some datasets, as well as differences between individuals (Supplementary Figures S11-S19). For example, in the Rattlesnake dataset, Move shows higher variance in raw accelerometer values compared to Not Moving, although to different degrees in different individuals.
Table 2: Summary of datasets in BEBE. Out of nine datasets, one comes from humans, three come from other terrestrial species, three come from aquatic species, and two come from flying species. Datasets marked with an asterisk are publicly available for the first time in BEBE. Duration: the total duration of the dataset. Duration annotated: the duration of annotated data. Mean annot. dur.: the average duration an individual spends in a behavioral state (described in Methods). Annot. method: the type of data that was used to make the behavioral annotations. $C$: Creative Commons.